# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v1

@@ Abstract
Label Smoothing (LS) is widely adopted to curb overconfidence in neural network predictions and enhance generalization. However, previous research shows that LS can force feature representations into excessively tight clusters, eroding intra-class distinctions. More recent findings suggest that LS also induces overconfidence in misclassifications, yet the precise mechanism remained unclear. In this work, we decompose the loss term introduced by LS, revealing two key components: (i) a regularization term that functions only when the prediction is correct, and (ii) an error-enhancement term that emerges under misclassifications. This latter term compels the model to reinforce incorrect predictions with exaggerated certainty, further collapsing the feature space. To address these issues, we propose Max Suppression (MaxSup), which uniformly applies the intended regularization to both correct and incorrect predictions by penalizing the top-1 logit instead of the ground-truth logit. Through feature analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries. Extensive experiments on image classification and downstream tasks confirm that MaxSup is a more robust alternative to LS.  [^1] Code is available at: https://github.com/ZhouYuxuanYX/Maximum-Suppression-Regularization.
Machine Learning, ICML
@@
1 Introduction
Multi-class classification (Russakovsky et al., 2015; LeCun, 1998) conventionally relies on one-hot labels, implicitly treating each class as if it were completely orthogonal to every other. In reality, however, classes often share low-level attributes (Zeiler & Fergus, 2014; Silla & Freitas, 2011) or exhibit high-level semantic similarities (Chen et al., 2021; Yi et al., 2022; Novack et al., 2023), which makes the strict orthogonality assumption overly simplistic. This mismatch can lead to over-confident classifiers, ultimately reducing generalization (Guo et al., 2020).
To address overconfidence, Szegedy et al. (2016) introduced Label Smoothing (LS), blending a uniform distribution with the hard label to reduce the model’s certainty in the target class. LS has become a mainstay in both image recognition (He et al., 2016; Touvron et al., 2021; Liu et al., 2021; Zhou et al., 2022b) and neural machine translation (Gao et al., 2020; Alves et al., 2023), often improving accuracy and calibration (Müller et al., 2019). However, studies have also revealed that LS can produce overly tight clusters in the feature space (Kornblith et al., 2021; Sariyildiz et al., 2022; Xu & Liu, 2023), thereby lowering intra-class diversity and harming transferability (Feng et al., 2021). Meanwhile, Zhu et al. (2022) reported that LS inadvertently increases confidence in incorrect predictions, though the exact cause remained unclear.
In this paper, we show that LS’s training objective inherently contains an error-enhancement term that amplifies misclassified predictions, thus causing overconfident errors and tighter feature clusters (Section 3.1, Table 1). Extending Zhu et al. (2022), we define “overconfidence” in terms of the network’s top-1 prediction rather than calibration-based criteria. Our analysis further demonstrates that penalizing the ground-truth logit in misclassifications compresses the feature space, reducing intra-class variation (Table 2), as corroborated by Grad-CAM visualizations (Figure 3).
To overcome these limitations, we propose Max Suppression (MaxSup), which retains LS’s desirable regularization effect while eliminating its error-enhancement component. Rather than penalizing the ground-truth logit, MaxSup penalizes the largest logit, thus providing consistent regularization regardless of prediction correctness. By preventing ground-truth suppression during misclassification, MaxSup preserves richer intra-class variation and improves inter-class separability. As illustrated in Figure 1, this alleviates the compression and attentional shortcomings introduced by LS, leading to more robust feature representations. Extensive experiments on both image classification and semantic segmentation confirm that MaxSup not only alleviates intra-class collapse but also boosts final accuracy, enhances generalization, and strengthens transfer performance.
Our contributions are summarized as follows:
  • We present a logit-level analysis of Label Smoothing that unearths an ‘error-enhancement” term, revealing how LS inadvertently reinforces overconfidence in misclassified samples.
  • We propose Max Suppression (MaxSup), which preserves LS’s desired regularization while removing its detrimental error-enhancement component, thereby reducing intra-class compression and boosting both classification and downstream task performance.
@@ 2 Related Work
We survey regularization techniques before focusing on Label Smoothing (LS) and highlighting how MaxSup differs.
@@ 2.1 Regularization
Regularization enhances the generalization of deep neural networks by limiting model complexity through various strategies. Classical approaches such as $\ell_{2}$ (Krogh & Hertz, 1991) and $\ell_{1}$ (Zou & Hastie, 2005) regularization constrain large or sparse weights, respectively, while Dropout (Srivastava et al., 2014) randomly deactivates neurons to prevent feature co-adaptation. Among loss-based methods, Label Smoothing (LS) (Szegedy et al., 2016) redistributes probability mass away from the ground-truth class, improving both accuracy and calibration (Müller et al., 2019). Variants like Online Label Smoothing (OLS) (Zhang et al., 2021) and Zipf Label Smoothing (Zipf-LS) (Liang et al., 2022) adapt LS by considering the model’s evolving predictions, yet they still fail to address the fundamental issue that arises when the ground-truth logit is not maximal (Section 3.1, Table 1). Other loss-based regularizers, such as Confidence Penalty (Pereyra et al., 2017) and Logit Penalty (Dauphin & Cubuk, 2021), target different aspects of the output distribution. Confidence Penalty discourages overconfident predictions, whereas Logit Penalty minimizes the global $\ell_{2}$-norm of logits to improve feature separability (Kornblith et al., 2021). However, Logit Penalty can reduce intra-class variation, impairing transfer learning (Section 4).
Our MaxSup approach
MaxSup diverges from these methods by selectively penalizing only the top-1 logit $(z_{\textit{max}})$ rather than the ground-truth logit $(z_{gt})$. Unlike LS-based techniques, which can exacerbate errors by excessively shrinking $z_{gt}$ for misclassified samples, MaxSup uniformly applies regularization to all predictions, regardless of correctness. Consequently, it effectively sidesteps the error-enhancement issue, preserves richer intra-class diversity (Table 2), and sustains robust transfer performance across various datasets and architectures (Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing (LS) has also been extensively examined in the context of knowledge distillation. Yuan et al. (2020) showed that LS can act as a proxy for distillation, while Shen et al. (2021) explored its role within teacher–student frameworks. Chandrasegaran et al. (2022) further demonstrated that a low-temperature, LS-trained teacher can improve distillation performance. Meanwhile, Kornblith et al. (2021) found that LS tightens class clusters in feature space, reducing transfer performance. From a Neural Collapse (NC) perspective (Zhou et al., 2022a; Guo et al., 2024), LS drives the model toward rigid feature clusters, a phenomenon measured by Xu & Liu (2023) via a variability metric.
Comparison with existing LS techniques
Our primary objective is to mitigate the error-enhancement effect. Instead of refining a smoothed label, as in OLS or Zipf-LS, MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This simple yet effective modification ensures uniform regularization even when $z_{gt}$ is not the top logit, thereby maintaining greater intra-class diversity and avoiding the performance degradation common to LS-based approaches (Section 3.2). Additionally, MaxSup integrates seamlessly with standard training pipelines, requiring no extra computational overhead beyond simply replacing LS.
@@ 3 Max Suppression Regularization (MaxSup)
We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS). By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error-enhancement term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance. To address these shortcomings, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than to $z_{gt}$ (Equation 8, Section 3.2). This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability. Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3). The following sections elaborate on MaxSup’s formulation and its integration into the overall training pipeline.
@@ 3.1 Revisiting Label Smoothing
For a classification task with $K$ classes, LS converts a one-hot label $\mathbf{y}\in\mathbb{R}^{K}$ into a soft label $\mathbf{s}\in\mathbb{R}^{K}$:
@@
Here, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS. This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting. (See Appendix A for a formal proof.)
@@
$L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl{(}z_{gt}-z_{m}\Bigr{)}}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl{(}z_{gt}-z_{n}\Bigr{)}}_{\text{Error-Enhancement}},$ (5)
where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error-enhancement term vanishes when $z_{gt}=z_{\textit{max}}$.
@@
    Error-Enhancement: Penalizes the gap between $z_{gt}$ and larger logits, inadvertently increasing overconfidence in incorrect predictions.
Although LS aims to combat overfitting by reducing prediction confidence, its error-enhancement component can be detrimental for misclassified samples, as it widens the gap between the ground-truth logit $z_{gt}$ and the incorrect top logit. Concretely:
@@
    Correct Predictions $(z_{gt}=z_{\textit{max}})$: The error-enhancement term is zero, and the regularization term effectively reduces overconfidence by shrinking the gap between $z_{gt}$ and any smaller logits.
@@
    • Error-Enhancement: Increases the gap between $z_{gt}$ and larger logits, reinforcing overconfidence in incorrect predictions.
@@
These issues with LS on misclassified samples have also been observed in prior work (Xia et al., 2024). By precisely identifying these two components (regularization vs. error-enhancement), we can design a more targeted solution.
Ablation Study on LS Components
To gauge the influence of each LS component, we conduct an ablation study on ImageNet-1K using a DeiT-Small model (Touvron et al., 2021) without Mixup or CutMix. As shown in Table 1, LS’s gains arise solely from the regularization term, whereas the error-enhancement term degrades performance. In contrast, our proposed Max Suppression Regularization (MaxSup) omits the error-enhancement effect and achieves higher accuracy by retaining the beneficial regularization.
Table 1: Ablation on LS components using DeiT-Small on ImageNet-1K (without CutMix or Mixup). “Regularization” denotes penalizing logits smaller than $z_{gt}$; “Error-Enhancement” penalizes logits larger than $z_{gt}$. MaxSup removes error-enhancement while retaining regularization.
@@
+ Error-Enhancement $\tfrac{\alpha}{N}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$ 73.63
+ Error-Enhancement $\alpha\,(z_{gt}-z_{\textit{max}})$ 73.69
@@
From Table 1, it is evident that LS’s overall accuracy boost is exclusively attributed to the regularization component, whereas error-enhancement consistently degrades performance (73.63% or 73.69%). Removing the error-enhancement term while keeping only the regularization improves accuracy slightly (75.98% vs. 75.91%). Finally, by avoiding error-enhancement entirely and preserving the helpful regularization, MaxSup achieves 76.12% accuracy—surpassing LS. This result underscores that MaxSup directly addresses LS’s primary shortcoming by consistently applying the intended regularization even when the model’s top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Label Smoothing (LS) suffers from two main limitations: inconsistent regularization and error amplification. As discussed in Section 3.1 and illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even for misclassified examples, thereby unnecessarily widening the gap between $z_{gt}$ and the incorrect top-1 logit. To address these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$. This crucial shift ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue seen in LS (Table 1), and preserving the integrity of the ground-truth logit for more stable and robust learning performance.
@@
Logit-Level Formulation of MaxSup
Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
@@
Comparison with Label Smoothing
MaxSup fundamentally differs from LS in handling correct and incorrect predictions. When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence. However, when $z_{gt}\neq z_{\textit{max}}$, LS continues to shrink $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression. As illustrated in Figure 3, this allows the model to recover from mistakes more effectively and avoid reinforcing incorrect predictions.
Gradient Analysis
To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
@@
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits increase slightly by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is therefore spared from penalization, thereby avoiding the error-amplification issue seen in LS. For completeness, Appendix A provides a full derivation of these gradients, and Figure 2 compares the resulting logit distributions under different regularizers.
Behavior Across Different Samples
@@
MaxSup applies a dynamic penalty that depends on the model’s current predictions. For high-confidence, correctly classified examples, it behaves similarly to LS by reducing overconfidence, thus effectively mitigating overfitting. In contrast, for misclassified or uncertain samples, MaxSup specifically suppresses the incorrect top-1 logit, further safeguarding the ground-truth logit $z_{gt}$. This selective strategy preserves an accurate representation of the true class while actively discouraging the propagation of errors. As shown in Section 5.1 and Table 5, this promotes more robust decision boundaries and ultimately leads to stronger generalization performance.
Theoretical Insights and Practical Benefits
@@
MaxSup provides both theoretical and practical advantages compared to LS. Whereas LS applies a uniform penalty to the ground-truth logit regardless of correctness, MaxSup focuses on penalizing only the most confident logit $z_{\textit{max}}$. This dynamic adjustment prevents error accumulation in misclassifications, thereby ensuring more stable convergence. As a result, MaxSup achieves stronger generalization, exhibits greater robustness to label noise, and performs well on challenging datasets. Moreover, as shown in Section 4, MaxSup preserves higher intra-class diversity, which substantially improves transfer learning performance (Table 3) and yields more interpretable activation maps (Figure 3).
4 Analysis of MaxSup’s Learning Benefits
MaxSup simultaneously promotes inter-class separability and intra-class variation, both essential for robust classification and effective feature transfer. In this section, we explore how MaxSup achieves these objectives and contrast its effectiveness with alternative regularization methods.
4.1 Intra-Class Variation and Transferability
As noted in Section 3.1, Label Smoothing (LS) primarily restricts overconfidence when the ground-truth class is correctly predicted, inadvertently causing error enhancement for misclassified samples. This selective penalty can overly compress intra-class diversity. In contrast, MaxSup uniformly penalizes the top-1 logit in both correct and incorrect cases, eliminating LS’s error-enhancement component and thus preserving more fine-grained distinctions within each class. Table 2 compares intra-class variation $\bar{d}_{\text{within}}$ and inter-class separability $R^{2}$ (Kornblith et al., 2021) for a ResNet-50 model trained on ImageNet-1K. Although all regularization strategies reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup shows the smallest reduction, implying stronger retention of within-class variability—often correlated with improved generalization and transferability. The benefits of this richer intra-class structure appear clearly in Table 3, where linear transfer performance on CIFAR-10 is reported. Although LS and Logit Penalty improve ImageNet accuracy, they diminish transfer accuracy by over-suppressing informative features. In contrast, MaxSup preserves transfer performance near that of the baseline, suggesting it retains crucial, discriminative features that generalize effectively to downstream tasks.

### v2

@@ Abstract
Label Smoothing (LS) is widely adopted to reduce overconfidence in neural network predictions and improve generalization. Despite these benefits, recent studies reveal two critical issues with LS. First, LS induces overconfidence in misclassified samples. Second, it compacts feature representations into overly tight clusters, diluting intra-class diversity, although the precise cause of this phenomenon remained elusive. In this paper, we analytically decompose the LS-induced loss, exposing two key terms: (i) a regularization term that dampens overconfidence only when the prediction is correct, and (ii) an error-amplification term that arises under misclassifications. This latter term compels the network to reinforce incorrect predictions with undue certainty, exacerbating representation collapse. To address these shortcomings, we propose Max Suppression (MaxSup), which applies uniform regularization to both correct and incorrect predictions by penalizing the top-1 logit rather than the ground-truth logit. Through extensive feature-space analyses, we show that MaxSup restores intra-class variation and sharpens inter-class boundaries. Experiments on large-scale image classification and multiple downstream tasks confirm that MaxSup is a more robust alternative to LS, consistently reducing overconfidence while preserving richer feature representations.  [^1] Code is available at: https://anonymous.4open.science/r/Maximum-Suppression-Regularization-DB0C.
1 Introduction
Multi-class classification [17, 24] typically relies on one-hot labels, which implicitly treat different classes as mutually orthogonal. In practice, however, classes often share low-level features [29, 40] or exhibit high-level semantic similarities [3, 22, 38], rendering the one-hot assumption overly restrictive. Such a mismatch can yield over-confident classifiers and ultimately degrade generalization [10].
To moderate overconfidence, Szegedy et al. [32] introduced Label Smoothing (LS), which combines a uniform distribution with the hard ground-truth label, thereby reducing the model’s certainty in the primary class. LS has since become prevalent in image recognition [11, 20, 33, 43] and neural machine translation [1, 6], often boosting accuracy and calibration [21]. Yet subsequent work indicates that LS can overly compress features into tight clusters [13, 26, 37], hindering intra-class variability and transferability [5]. In parallel, Zhu et al. [44] found that LS paradoxically fosters overconfidence in misclassified samples, though the precise mechanism behind this remains uncertain.
In this paper, we reveal that LS’s training objective inherently contains an error amplification term. This term pushes the network to reinforce incorrect predictions with exaggerated certainty, yielding highly confident misclassifications and further compressing feature clusters (Section 3.1, Table 1). Building on Zhu et al. [44], we characterize “overconfidence” in terms of the model’s top-1 prediction, rather than through conventional calibration metrics. Through our analysis, we further show that punishing the ground-truth logit during misclassification reduces intra-class variation (Table 2), a phenomenon corroborated by Grad-CAM visualizations (Figure 2).
To overcome these shortcomings, we introduce Max Suppression (MaxSup), a method that retains the beneficial regularization effect of LS while eliminating its error amplification. Rather than penalizing the ground-truth logit, MaxSup focuses on the model’s top-1 logit, ensuring a consistent regularization signal regardless of whether the current prediction is correct or misclassified. By preserving the ground-truth logit in misclassifications, MaxSup sustains richer intra-class variability and sharpens inter-class boundaries. As visualized in Figure 1, this approach mitigates the feature collapse and attention drift often induced by LS, ultimately leading to more robust representations. Through comprehensive experiments in both image classification (Section 4.2) and semantic segmentation (Section 4.3), we show that MaxSup not only alleviates severe intra-class collapse but also consistently boosts top-1 accuracy (+0.4%–1.0% improvement over LS on ImageNet) and robustly enhances downstream transfer performance (Section 4.1).
Our contributions are summarized as follows:
  • We perform a logit-level analysis of Label Smoothing, revealing how the error amplification term inflates misclassification confidence and compresses features.
  • We propose Max Suppression (MaxSup), removing detrimental error amplification while preserving LS’s beneficial regularization. As shown in extensive ablations, MaxSup alleviates intra-class collapse and yields consistent accuracy gains.
  • We demonstrate superior performance across tasks and architectures, including ResNet, MobileNetV2, and DeiT-S, where MaxSup significantly boosts accuracy on ImageNet (up to +1.0% in top-1) and consistently delivers stronger representations for downstream tasks such as semantic segmentation and robust transfer learning.
@@ 2 Related Work
We first outline mainstream regularization techniques in deep learning, then survey recent advances in Label Smoothing (LS), and finally clarify how our MaxSup diverges from prior variants.
@@ 2.1 Regularization
Regularization techniques aim to improve the generalization of deep neural networks by constraining model complexity. Classical methods like $\ell_{2}$ [16] and $\ell_{1}$ [45] impose direct penalties on large or sparse weights, while Dropout [30] randomly deactivates neurons to discourage over-adaptation. In the realm of loss-based strategies, Label Smoothing (LS) [32] redistributes a fraction of the label probability mass away from the ground-truth class, thereby improving accuracy and calibration [21]. Variants such as Online Label Smoothing (OLS) [41] and Zipf Label Smoothing (Zipf-LS) [19] refine LS by dynamically adjusting the smoothed labels based on a model’s evolving predictions. However, they do not fully address the fundamental issue that emerges when the ground-truth logit is not the highest one (see Section 3.1, Table 1). Other loss-based regularizers focus on alternative aspects of the predictive distribution. Confidence Penalty [23] penalizes the model’s confidence directly, while Logit Penalty [4] minimizes the global $\ell_{2}$-norm of logits, a technique reported to enhance class separation [13]. Despite these benefits, Logit Penalty can inadvertently shrink intra-class variation, thereby hampering transfer learning (see Section 4.1). Unlike the aforementioned methods, MaxSup enforces regularization by penalizing only the top-1 logit $z_{\textit{max}}$ rather than the ground-truth logit $z_{gt}$. In LS-based approaches, suppressing $z_{gt}$ for misclassified samples can worsen errors, whereas MaxSup applies a uniform penalty regardless of whether the model’s prediction is correct. Consequently, MaxSup avoids the error amplification effect, retains richer intra-class diversity (see Table 2), and achieves robust transfer performance across diverse datasets and model families (see Table 3).
@@ 2.2 Studies on Label Smoothing
Label Smoothing has also been studied extensively under knowledge distillation. For instance, Yuan et al. [39] observed that LS can approximate the effect of a teacher–student framework, while Shen et al. [28] investigated its role in such pipelines more systematically. Additionally, Chandrasegaran et al. [2] demonstrated that a low-temperature, LS-trained teacher can notably improve distillation outcomes. Concurrently, Kornblith et al. [13] showed that LS tightens intra-class clusters in the feature space, diminishing transfer performance. From a Neural Collapse perspective [42, 9], LS nudges the model toward rigid feature clusters, as evidenced by the reduced feature variability measured in Xu and Liu [37]. Our goal is to overcome LS’s inherent error amplification effect. Rather than adjusting how the smoothed label distribution is constructed (as in OLS or Zipf-LS), MaxSup directly penalizes the highest logit $z_{\textit{max}}$. This design ensures consistent regularization even if $z_{gt}$ is not the top logit, thereby avoiding the degradation in performance typical of misclassified samples under LS (see Section 3.2). Moreover, MaxSup integrates seamlessly into standard training pipelines, introducing negligible computational overhead beyond substituting the LS term.
@@ 3 Max Suppression Regularization (MaxSup)
We first partition the training objective into two components: the standard Cross-Entropy (CE) loss and a regularization term introduced by Label Smoothing (LS). By expressing LS in terms of logits (Theorem 3.3), we isolate two key factors: a regularization term that controls overconfidence and an error amplification term that enlarges the gap between the ground-truth logit $z_{gt}$ and any higher logits (Corollary 3.4, Equation 5), ultimately degrading performance. To address these issues, we propose Max Suppression Regularization (MaxSup), which applies the penalty to the largest logit $z_{\textit{max}}$ rather than $z_{gt}$ (Equation 8, Section 3.2). This shift delivers consistent regularization for both correct and incorrect predictions, preserves intra-class variation, and bolsters inter-class separability. Consequently, MaxSup mitigates the representation collapse found in LS, attains superior ImageNet-1K accuracy (Table 1), and improves transferability (Table 2, Table 3). The following sections elaborate on MaxSup’s formulation and integration into the training pipeline.
@@ 3.1 Revisiting Label Smoothing
For a standard classification task with $K$ classes, Label Smoothing (LS) converts a one-hot label $\mathbf{y}\in\mathbb{R}^{K}$ into a softened target label $\mathbf{s}\in\mathbb{R}^{K}$:
@@
Where, $\mathbf{q}$ is the predicted probability vector, $H(\cdot)$ denotes the Cross-Entropy, and $\frac{\mathbf{1}}{K}$ is the uniform distribution introduced by LS. This shows that LS adds a regularization term, $L_{\textit{LS}}$, which smooths the output distribution and helps to reduce overfitting. (See Appendix A for a formal proof.)
@@
$L_{\textit{LS}}\;=\;\underbrace{\frac{\alpha}{K}\sum_{z_{m}<z_{gt}}\Bigl{(}z_{gt}-z_{m}\Bigr{)}}_{\text{Regularization}}\;+\;\underbrace{\frac{\alpha}{K}\sum_{z_{n}>z_{gt}}\Bigl{(}z_{gt}-z_{n}\Bigr{)}}_{\text{error amplification}},$ (5)
where $M$ and $N$ are the numbers of logits below and above $z_{gt}$, respectively ($M+N=K-1$). Note that the error amplification term vanishes when $z_{gt}=z_{\textit{max}}$.
@@
    Error amplification: Penalizes the gap between $z_{gt}$ and larger logits, inadvertently increasing overconfidence in incorrect predictions.
Although LS aims to combat overfitting by reducing prediction confidence, its error amplification component can be detrimental for misclassified samples, as it widens the gap between the ground-truth logit $z_{gt}$ and the incorrect top logit. Concretely:
@@
    Correct Predictions $(z_{gt}=z_{\textit{max}})$: The error amplification term is zero, and the regularization term effectively reduces overconfidence by shrinking the gap between $z_{gt}$ and any smaller logits.
@@
    • Error amplification: Increases the gap between $z_{gt}$ and larger logits, reinforcing overconfidence in incorrect predictions.
@@
These issues with LS on misclassified samples have also been systematically observed in prior work [35]. By precisely disentangling these two components (regularization vs. error amplification), we can design a more targeted and effective solution.
Table 1: Ablation on LS components using DeiT-Small on ImageNet-1K (without CutMix or Mixup). “Regularization” denotes penalizing logits smaller than $z_{gt}$; “error amplification” penalizes logits larger than $z_{gt}$. MaxSup removes error amplification while retaining regularization.
@@
+ error amplification $\tfrac{\alpha}{N}\sum_{z_{n}>z_{gt}}(z_{gt}-z_{n})$ 73.63
+ error amplification $\alpha\,(z_{gt}-z_{\textit{max}})$ 73.69
@@
Ablation Study on LS Components. To isolate the effects of each component in LS, we carefully perform a detailed and systematic ablation study on ImageNet-1K using a DeiT-Small model [33] without Mixup or CutMix. As indicated in Table 1, the performance gains from LS stem solely from the regularization term, whereas the error amplification term degrades accuracy. In contrast, our MaxSup omits the error amplification component and leverages only the beneficial regularization, thereby boosting accuracy beyond that of standard LS. Specifically, Table 1 shows that LS’s overall improvement can be attributed exclusively to its regularization contribution; the error amplification term consistently reduces accuracy (e.g., to 73.63% or 73.69%). Disabling only the error amplification while retaining the regularization yields a slight but measurable improvement (75.98% vs. 75.91%). By fully removing error amplification and faithfully preserving the helpful aspects of LS, our MaxSup achieves 76.12% accuracy, clearly and consistently outperforming LS. This result underscores that MaxSup directly tackles LS’s fundamental shortcoming by maintaining a consistent and meaningful regularization signal—even when the top-1 prediction is incorrect.
@@ 3.2 Max Suppression Regularization
Building on our analysis in Section 3.1, we find that Label Smoothing (LS) not only impacts correctly classified samples but also influences misclassifications in unintended and harmful ways. Specifically, LS suffers from two main limitations: inconsistent regularization and error amplification. As illustrated in Table 1, LS penalizes the ground-truth logit $z_{gt}$ even in misclassified examples, needlessly widening the gap between $z_{gt}$ and the erroneous top-1 logit. To resolve these critical shortcomings, we propose Max Suppression Regularization (MaxSup), which explicitly penalizes the largest logit $z_{\textit{max}}$ rather than $z_{gt}$. This key design choice ensures uniform regularization across both correct and misclassified samples, effectively eliminating the error-amplification issue in LS (Table 1) and preserving the ground-truth logit’s integrity for more stable, robust learning.
@@
Logit-Level Formulation of MaxSup. Building on the logit-level perspective introduced for LS in Section 3.1, we can express $L_{\textit{MaxSup}}$ as:
@@
Comparison with Label Smoothing. MaxSup fundamentally differs from LS in handling correct and incorrect predictions. When $z_{gt}=z_{\textit{max}}$, both LS and MaxSup similarly reduce overconfidence. However, when $z_{gt}\neq z_{\textit{max}}$, LS shrinks $z_{gt}$, widening the gap with the incorrect logit, whereas MaxSup penalizes $z_{\textit{max}}$, preserving $z_{gt}$ from undue suppression. As illustrated in Figure 2, this helps the model recover from mistakes more effectively and avoid reinforcing incorrect predictions.
Gradient Analysis. To understand MaxSup’s optimization dynamics, we compute its gradients with respect to each logit $z_{k}$. Specifically,
@@
Thus, the top-1 logit $z_{\textit{max}}$ is reduced by $\alpha\bigl{(}1-\tfrac{1}{K}\bigr{)}$, while all other logits slightly increase by $\tfrac{\alpha}{K}$. In misclassified cases, the ground-truth logit $z_{gt}$ is spared from penalization, avoiding the error-amplification issue seen in LS. For completeness, Appendix A provides the full gradient derivation.
@@
Behavior Across Different Samples. MaxSup applies a dynamic penalty based on the model’s current predictions. For high-confidence, correctly classified examples, it behaves similarly to LS by reducing overconfidence, effectively mitigating overfitting. In contrast, for misclassified or uncertain samples, MaxSup aggressively suppresses the incorrect top-1 logit, further safeguarding the ground-truth logit $z_{gt}$. This selective strategy preserves a faithful and reliable representation of the true class while actively discouraging error propagation. As shown in Section 4.2 and Table 5, this promotes more robust decision boundaries and leads to stronger generalization.
@@
Theoretical Insights and Practical Benefits. MaxSup provides both theoretical and practical advantages over LS. Whereas LS applies a uniform penalty to the ground-truth logit regardless of correctness, MaxSup penalizes only the most confident logit $z_{\textit{max}}$. This dynamic adjustment robustly prevents error accumulation in misclassifications, ensuring more stable convergence. As a result, MaxSup generalizes better and achieves strong performance on challenging datasets. Moreover, as shown in Section 4.1, MaxSup preserves greater intra-class diversity, substantially improving transfer learning (Table 3) and yielding more interpretable activation maps (Figure 2).
4 Experiments
We begin by examining how MaxSup improves feature representations, then evaluate it on large-scale image classification and semantic segmentation tasks. Finally, we visualize class activation maps to illustrate the practical benefits of MaxSup.
4.1 Analysis of MaxSup’s Learning Benefits
Having established how MaxSup addresses Label Smoothing’s (LS) principal shortcomings (Section 3.1), we now demonstrate its impact on inter-class separability and intra-class variation—two properties essential for accurate classification and effective transfer learning.